OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x - sin x

OpenStudy (anonymous):

Answers are: A. cos^2x-sin^2x-sinx B. cos^sx-sin^3x C. cos^sx+sin^2x+sinx D. cos^sx-3sinx

OpenStudy (anonymous):

cos(a+b)=cosa cos b -sin a sin b put a=b=x cos(x+x)=cos x cos x-sin x sin x \[\cos 2x=\cos ^{2}x-\sin ^{2}x\] now you can write the solution

OpenStudy (anonymous):

how do i do that? thanks btw :)

OpenStudy (anonymous):

\[\cos 2x-\sin x=\cos ^{2}x-\sin ^{2}x-\sin x\]

OpenStudy (anonymous):

Thank yu so mucH!!!

OpenStudy (anonymous):

yw